Miller Company’s contribution format income statement for the most recent month is shown below:

Total Per Unit
Sales (41,000 units) $ 246,000 $ 6.00
Variable expenses 123,000 3.00
Contribution margin 123,000 $ 3.00
Fixed expenses 50,000
Net operating income $ 73,000

Required:

(Consider each case independently):

1. What is the revised net operating income if unit sales increase by 11%?

2. What is the revised net operating income if the selling price decreases by $1.20 per unit and the number of units sold increases by 17%?

3. What is the revised net operating income if the selling price increases by $1.20 per unit, fixed expenses increase by $7,000, and the number of units sold decreases by 7%?

4. What is the revised net operating income if the selling price per unit increases by 20%, variable expenses increase by 30 cents per unit, and the number of units sold decreases by 10%?
Answer #1

Ans:

1

Net operating income

$86,530

2

Net operating income

$36,346

3

Net operating income

$103,146

4

Net operating income

$93,910

Working:

1) If unit sale increased by 11%

41,000units*111%=45,510units

Per unit $

Units

Total $

Revenue

         6.00

       45,510

         273,060

Variable expense

       (3.00)

       45,510

       (136,530)

Contribution margin

         3.00

        136,530

Fixed cost

         (50,000)

Net Operating Income

           $86,530

2) If sale price decreases by $1.20per unit and sale unit increase by 17%

Revised selling price = 6-1.2=$4.8

Revised sales units = 41,000units*117%=47,970units

Per unit $

Units

Total $

Revenue

         4.80

       47,970

         230,256

Variable expense

       (3.00)

       47,970

       (143,910)

Contribution margin

         1.80

           86,346

Fixed cost

         (50,000)

Net Operating Income

           $36,346

3) If sale price increase by $1.20 per unit, fixed expense increase by $7,000 and sale unit decrease by 7%

Revised selling price = 6+1.2=$7.2

Revised sales units = 41,000units*93%=38,130units

Per unit $

Units

Total $

Revenue

         7.20

       38,130

         274,536

Variable expense

       (3.00)

       38,130

      (114,390)

Contribution margin

         4.20

        160,146

Fixed cost (50,000+7,000)

         (57,000)

Net Operating Income

           $103,146

4) If sale price increase by 20%, variable expense increase by 30 cents and sale unit decrease by 10%

Revised selling price = $6*120%=$7.2

Revised sales units = 41,000units*90%=36,900units

Per unit $

Units

Total $

Revenue

         7.20

       36,900

         265,680

Variable expense $3+30cents=$3.30

       (3.30)

       36,900

       (121,770)

Contribution margin

3.90

        143,910

Fixed cost

         (50,000)

Net Operating Income

           $93,910